OpenStudy (anonymous):

-Under the Quadratic equation (ROOTS) Determine the possible values of k so that the roots of 4x^2+kx+35=0 differ by 1

OpenStudy (anonymous):

let the two roots be alpha and beta alhpha + beta = -k/4 (alpha)(beta)=35/4 (alpha + beta ) ^2 = ( alpha - beta )^2 + 4(alpha)(beta) k^2/16 = 1+35 k^2 = 36(16) k = 6* 4 = 24